OpenStudy (anonymous):

Find the product. (-2x^2)^3 ·3x

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) The first step is multiplying everything within the brackets to the power of three. \[(-2x^2)^2 = 2^3x ^{2 * 3}\]

OpenStudy (loser66):

mistake, @Compassionate , recheck , please

OpenStudy (compassionate):

Negative 2.

OpenStudy (loser66):

-24, you miss the last 3

OpenStudy (loser66):

ok?

OpenStudy (anonymous):